OpenStudy (anonymous):

convert the rectangular equation to polar form. assume a is less than 0. x=10

OpenStudy (amistre64):

(10,0)

OpenStudy (amistre64):

oh, the equation, not the point lol

OpenStudy (anonymous):

yeah haha could you please show the work? thank you

OpenStudy (amistre64):

r cos(t) =10 r = 10/cos(t)

OpenStudy (anonymous):

where did u get t from?

OpenStudy (amistre64):

well, x = r cos(t) thats just by definition

OpenStudy (amistre64):

so, r cos(t) = 10 is a direct translation to polar form

OpenStudy (anonymous):

do you mean theta?

OpenStudy (amistre64):

i dont think I did the divide right tho

OpenStudy (amistre64):

yes; t = theta ... it really doesnt matter what you call it, it never comes when called

OpenStudy (anonymous):

ohh ok thanks i was just checking because i was confused. anyways so the answer is r=10/cos theta

OpenStudy (amistre64):

maybe, i cant remember if the divide by 0 applies to mess things up with

OpenStudy (anonymous):

im not sure either its quite confusing

OpenStudy (amistre64):

I know to get a straight line; which is what x=10 is; you keep the theta a constant... but id have to read up on it again

OpenStudy (anonymous):

i think tthe answer is right because some of my other answers ended with r=///// so it should be right hold on no the answer is r=10 sec theta

OpenStudy (anonymous):

how do you get secant?

OpenStudy (amistre64):

r = sec and 1/cos are the same thing

OpenStudy (amistre64):

\[sec = \frac{1}{cos}\]

OpenStudy (amistre64):

\[\frac{10}{cos}=10 sec\]

OpenStudy (anonymous):

i see thanks so much for your explanations

OpenStudy (amistre64):

youre welcome :)